# Cohomogeneity one $\mathsf{RCD}$-spaces

### v2

@@ 2.1. Calculus on metric measure spaces
[…]
@@
Let $(X,d,\mathfrak{m})$ be a metric measure space. We define the Sobolev space $W^{1,2}(X):=\{f\in L^{2}(\mathfrak{m})\,|\,\mathrm{Ch}(f)<\infty\}$, equipped with the norm $\|f\|^{2}_{W^{1,2}}:=\|f\|^{2}_{2}+\mathrm{Ch}^{2}(f)$. In the case that this norm comes from an inner product, we say that $(X,d,\mathfrak{m})$ is infinitesimally Hilbertian.
[…]
@@
$\langle\nabla f,\nabla g\rangle:=\frac{1}{2}\left(|\nabla(f+g)|^{2}-|\nabla(f-g)|^{2}\right)\in L^{1}(\mathfrak{m}).$
Let us recall the following characterization of infinitesimal Hilbertianity (see Theorem $4.3.3$ in [36]).
@@
[…]
We denote the function $g$ (which if exists, is unique) by $\Delta f$ and refer to it as the Laplacian of $f$.
In case that the underlying space is infinitesimally Hilbertian we have at our disposal the following good properties of the Laplacian and of the pointwise inner product (see Theorem $4.3.3$ part $v),$ Remark $5.2.2,$ and Proposition $5.2.3$ in [36]).
@@
[…]
@@ 2.2. Bakry-Émery condition
[…]
@@
[…]
This fact was shown by Ambrosio, Gigli and Savaré for $N=\infty$ (under equivalent technical assumptions) in [3, Theorem 1.1] and for $N<\infty$ by Erbar, Kuwada and Sturm in [24, Theorem 7]. Note also the equivalence between the $\mathsf{CD}^{\ast}(K,N)$- and $\mathsf{CD}(K,N)$- conditions due to Cavalletti and Milman, [17, Corollary 13.7].
Since we only consider compact spaces with a Borel measure, it is clear that condition (2) above is automatically fulfilled.
From these considerations, we point out that we do not need to make use of the optimal transport formulation of the curvature-dimension condition, for example as stated in [40], so instead we settle on the following:
@@
One important thing to mention is that in general the Bakry-Émery condition is strictly weaker than the $\mathsf{RCD}$-condition. In [47] Honda gives examples of spaces that satisfy a $\mathsf{BE}(K,N)$ condition but that are not $\mathsf{RCD}$-spaces. The reason being that they do not satisfy the Sobolev-to-Lipschitz property. Moreover, the Sobolev-to-Lipschitz property is essential as the $\mathsf{BE}(K,N)$-condition alone cannot imply anything about the geometry of the space as was pointed out to us by Gigli with the following examples. Any metric space equipped with an atomic measure satisfies $\mathsf{BE}(K,N)$ for any $K$ and $N$ as the Cheeger energy is identically 00. A more geometric example is the following: For a given compact smooth embedded submanifold $M$ in $\mathbb{R}^{n}$ one can consider the induced Riemannian metric $g$ from $\mathbb{R}^{n}$ (and in turn the Riemannian distance $d_{g}$) and Riemannian volume $d\mathrm{vol}_{g}$. Then by compactness, the metric measure space $(M,d_{g},d\mathrm{vol}_{g})$ is an $\mathsf{RCD}(K,N)$-space for adequate $K$ and $N$. Now, $M$ can be endowed with the restriction of the Euclidean distance as well, i.e., $d(x,y)=\left\|x-y\right\|$ (and the same measure $d\mathrm{vol}_{g}$). It is clear that this distance in general does not make $M$ a geodesic space and thus, this metric measure structure cannot satisfy the $\mathsf{RCD}$-condition in general. However, the Cheeger energies associated to both distances agree as can be seen directly from the definition above since the Lipschitz constants of functions almost agree for both distances since they almost coincide at small scales.
Often, if there is no room for confusion, we simply refer to metric measure spaces satisfying Definition 2.9 as $\mathsf{RCD}$-spaces without making any explicit mention of the parameters $K$ and $N$.
@@
Let $(X,d,\mathfrak{m})\in\mathbf{X}$ be a length space and assume that there exists a covering $\{U_{i}\}_{i\in I}$ of $X$ by non-empty open subsets such that $\mathfrak{m}(U_{i})<\infty$ if $K<0$, and $(\overline{U}_{i},d,\mathfrak{m}\llcorner_{\overline{U}_{i}})\in X$ satisfy $\mathsf{RCD}(K,N)$. Then $(X,d,\mathfrak{m})$ is a $\mathsf{RCD}(K,N)$ space.
Let us describe some features on $\mathsf{RCD}$-spaces that will be useful to us. Proofs of them can be found for example in [24]. We remark that the tensorization property in item (2) below was originally shown to hold for the $\mathsf{BE}$-condition in [3].
@@
Let $(X,d,x),(Y,d_{Y},y)$ be complete and separable pointed metric spaces. Given $\epsilon>0$. We say that a map $f_{\epsilon}\colon B_{1/\epsilon}(x)\rightarrow B_{1/\epsilon}(y)$ is an $\epsilon$-Gromov-Hausdorff approximation if:
@@
    For all $u,v\in B_{1/\epsilon}(x)$ $|d(u,v)-d_{Y}(f_{\epsilon}(u),f_{\epsilon}(v))|<\epsilon$.
@@
    for all $y\in B_{1/\epsilon}(y)$ there exists some $x\in B_{1/\epsilon}(x)$ such that $d_{Y}(f_{\epsilon}(x),y)<\epsilon.$
@@
A sequence $\{(X_{n},d_{n},\mathfrak{m}_{n},x_{n})\}_{n\in\mathbb{N}}$ of complete pointed metric measure spaces converges in the pointed measured Gromov-Hausdorff sense to a complete pointed metric measure space $(Y,d_{Y},\mathfrak{m}_{Y},y)$ if it converges in the pointed Gromov-Hausdorff sense with the extra assumption that the $\epsilon_{n}$-Gromov-Hausdorff approximations are measurable and satisfy:
$f_{\epsilon\#}\mathfrak{m}_{n}\rightharpoonup\mathfrak{m}_{Y}.$
@@
Let $(X,d,\mathfrak{m})$ be an $\mathsf{RCD}(K,N)$-space for $K\in\mathbb{R}$ and $N\in[1,\infty)$. Given a point $x_{0}\in X$, we say that a pointed metric measure space $(Y,d_{Y},\mathfrak{m}_{Y},y)$ is a tangent of $X$ at $x_{0}$ if there exists a sequence $\{r_{i}\}_{i\in\mathbb{N}}\subset(0,1)$ such that $r_{i}\rightarrow 0$, and the sequence
@@
Given $m\in\mathbb{N}\cap[1,N]$, let $\mathcal{R}_{m}$ denote the set of points $x\in X$ such that $\mathrm{Tan}(X,d,\mathfrak{m},x)=\{(\mathbb{R}^{m},d_{\mathbb{E}},\mathcal{L}^{m},0)\}.$ Using regular Lagrangian flows, Brué and Semola in [12] proved that there is precisely a unique $m$ such that the set $\mathcal{R}_{m}$ has full measure:
@@ 2.3. Non-collapsed spaces
Thanks to the structure of $\mathsf{RCD}$-spaces, we can check that for non-collapsed spaces we have necessarily that $N\in\mathbb{N}$ (see Theorem $1.12$ in [23]). Non-collapsed $\mathsf{RCD}$-spaces also enjoy good structural properties which are stronger than those of more general $\mathsf{RCD}$-spaces. For example, we have the following result about the tangents (see Proposition $2.8$ in [23]):
@@
Regarding the essential dimension, we have the following result by Brena, Gigli, Honda, and Zhu [11, Theorem 1.7] which lets us identify non-collapsed spaces.
@@ 2.4. Warped products
[…]
Given an AC curve $\gamma\colon[0,1]\to X$, for a.e. $t_{0}\in[0,1]$ fixed the following limit exits
$|\dot{\gamma}(t_{0})|:=\lim_{h\to 0}\frac{d_{X}(\gamma(t_{0}+h),\gamma(t_{0})}{|h|}.$
Moreover the function $t\mapsto|\dot{\gamma}(t)|$ called metric speed of $\gamma$ is in $L^{1}(0,1)$, and is the minimal function satisfying (3).
@@
[…]
For convenience we define the length of a curve that is not as admissible simply as +.$\infty$.+ ∞ . $L_{f}$ is a length-structure on the class of admissible curves. We define a pseudo-metric $\tilde{d}_{f}$ on $B\times F$, between $(p,x)$ and $(q,y)$ by setting
@@
Then we have an induced equivalence relation $\sim$, where $(p,x)\sim(q,y)$ if and only if $\tilde{d}_{f}((p,x),(q,y))=0$. We define the warped product of $B$ and $F$ via $f$ as $B\times_{f}F:=B\times F/\sim$ with the distance $d_{f}([p,x],[q,y]):=\tilde{d}_{f}((p,x),(q,y))$. By definition $B\times_{f}F$ is an intrinsic metric space, and since $B$ and $F$ are complete and locally compact, then the warped product is complete and locally compact. As pointed out in [51, Remark 2.4], the warped product $B\times_{f}F$ is geodesic.
Moreover, as pointed out in [1] we have the following description of geodesics
Proposition 2.23 (Propsition 4.5 (2)).
Let $\gamma=(\alpha,\beta)$ be a geodesic in $B\times_{f}F$ such that $f\circ\alpha>0$. Then $\beta$ is a pregodesic in $F$. That is, up to a reparametrization $\beta$ is a geodesic in $F$.
Lemma 2.24.
Let $F$ be a compact Riemannian manifold. Consider two functions
$f,g\colon[0,\infty)\to F$
such that $f(0)=0=g(0)$, and for $t>0$ we have $f(t),g(t)>0$. Consider $0<s_{1}<s_{2}$, and $p,q\in F$ are such that there exists a unique geodesic between them. Then the trace of the geodesic $\gamma_{f}$ from $(s_{1},p)$ to $(s_{2},q)$ in $[0,\infty)\times_{f}F$ is the same as the trace of the geodesic $\gamma_{g}$ from $(s_{1},p)$ to $(s_{2},q)$ in $[0,\infty)\times_{g}F$.
Proof.
We write the geodesic $\gamma_{f}=(\alpha_{f},\beta_{f})$, and observe that by Proposition 2.23 $\beta_{f}$ is a pregeodesic in $F$. The same holds for the geodesic $\gamma_{g}=(\alpha_{g},\beta_{g})$ in $[0,\infty)\times_{g}F$. Thus $\beta_{f}$ and $\beta_{g}$ are two pregeodesics in $F$ joining $p$ to $q$. Since $p$ and $q$ are not conjugate this implies that the trace of $\beta_{f}$ is the same as the trace of $\beta_{g}$. ∎
@@
$\mathrm{Con}_{K}(F)=\begin{cases}F\times[0,\pi/\sqrt{K}]/(F\times\{0,\pi/\sqrt{K}\})&\mbox{if }K>0,\\ F\times[0,\infty)/(F\times\{0\})&\mbox{if }K\leq 0.\end{cases}$
@@
:=$:=\begin{cases}\cos^{-1}_{K}\Big{(}\cos_{K}(t)\cos_{K}(t)+K\sin_{K}(s)\sin_{K}(t)\cos_{K}\big{(}d_{F}(p,q)\wedge\pi\big{)}\Big{)}&\mbox{if }K\neq 0,\\ \sqrt{s^{2}+t^{2}-2st\cos\big{(}d_{F}(p,q)\wedge\pi\big{)}}&\mbox{if }K=0,\end{cases}$
@@
Here, for $I_{K}=[0,\pi/\sqrt{K}]$ for $K>0$ and $[0,\infty)$ for $K\leq 0$, we have $\cos_{K}\colon I_{K}\to[0,\infty)$ given by
$\cos_{K}(t):=\begin{cases}\cos\left(\sqrt{K}t\right)&\mbox{if }K>0,\\ \cosh\left(\sqrt{-K}t\right)&\mbox{if }K<0.\end{cases}$
@@
$\sin_{K}(t):=\begin{cases}\frac{1}{\sqrt{K}}\sin\left(\sqrt{K}t\right)&\mbox{if }K>0,\\ t&K=0,\\ \frac{1}{\sqrt{|K|}}\sinh\left(\sqrt{|K|}t\right)&\mbox{if }K<0.\\ \end{cases}$
@@
If $\operatorname{diam}(F)\leq\pi$ and $F$ is a length space, then the $(K,N)$-cone $\mathrm{Con}^{N}_{K}(F)$ coincides with the $N$-warped product $I_{K}\times_{\sin_{K}}^{N}F$.

### v3

@@ 2.1. Calculus on metric measure spaces
$\mathrm{Ch}(f):=\inf\left\{\liminf_{n\rightarrow\infty}\frac{1}{2}\int|\mathrm{Lip}f_{n}|^{2}d\mathfrak{m}\,|\,f_{n}\in\mathrm{LIP}(X),f_{n}\rightarrow f\text{ in }L^{2}(\mathfrak{m})\right\}.$
@@
Let $(X,d,\mathfrak{m})$ be a metric measure space. We define the Sobolev space $W^{1,2}(X):=\{f\in L^{2}(\mathfrak{m})\,|\,\mathrm{Ch}(f)<\infty\}$, equipped with the norm $\|f\|^{2}_{W^{1,2}}:=\|f\|^{2}_{2}+2\mathrm{Ch}(f)$. In the case that this norm comes from an inner product, we say that $(X,d,\mathfrak{m})$ is infinitesimally Hilbertian.
Given a function $f\in W^{1,2}(X)$, there exists a distinguished function $|\nabla f|\in L^{2}(\mathfrak{m})$ called the minimal weak upper gradient of $f$, which satisfies $\mathrm{Ch}(f)=\frac{1}{2}\int|\nabla f|^{2}d\mathfrak{m}$ (see Proposition $2.2.8$ in [34]).
@@
$\langle\nabla f,\nabla g\rangle:=\frac{1}{4}\left(|\nabla(f+g)|^{2}-|\nabla(f-g)|^{2}\right)\in L^{1}(\mathfrak{m}).$
Let us recall the following characterization of infinitesimal Hilbertianity (see Theorem 4.3.3. in [34]).
@@
$\int hg\,dm=-\int\langle\nabla f,\nabla h\rangle d\mathfrak{m},\quad\forall h\in W^{1,2}(X).$
We denote the function $g$ (which is unique if exists) by $\Delta f$ and refer to it as the Laplacian of $f$.
In the case where the underlying space is infinitesimally Hilbertian we have at our disposal the following good properties of the Laplacian and of the pointwise inner product (see Theorem 4.3.3 part $v)$, Remark 5.2.2, and Proposition 5.2.3 in [34]).
@@
  • The pointwise inner product $\langle\nabla\cdot,\nabla\cdot\rangle$ is continuous.
@@ 2.2. Bakry-Émery condition
(2) $\frac{1}{2}\int\Delta g|\nabla f|^{2}d\mathfrak{m}-\int g\langle\nabla(\Delta f),\nabla f\rangle d\mathfrak{m}\geq K\int g|\nabla f|^{2}d\mathfrak{m}+\frac{1}{N}\int g(\Delta f)^{2}d\mathfrak{m}.$
@@
    $\int\exp(-cd^{2}(x_{0},x))d\mathfrak{m}(x)<\infty.$
This fact was shown by Ambrosio, Gigli and Savaré for $N=\infty$ (under equivalent technical assumptions) in [6, Theorem 1.1] and for $N<\infty$ by Erbar, Kuwada and Sturm in [23, Theorem 7]. Note also the equivalence between the $\mathsf{CD}^{\ast}(K,N)$- and $\mathsf{CD}(K,N)$- conditions due to Cavalletti and Milman, [16, Corollary 13.7] and Li [56].
Since we only consider compact spaces with a Radon measure, it is clear that condition (2) above is automatically fulfilled.
From these considerations, we point out that we do not need to make use of the optimal transport formulation of the curvature-dimension condition, for example as stated in [38], so instead we settle on the following definition.
@@
One important thing to mention is that in general the Bakry-Émery condition is strictly weaker than the $\mathsf{RCD}$-condition. In [46] Honda gives examples of spaces that satisfy a $\mathsf{BE}(K,N)$ condition but that are not $\mathsf{RCD}$-spaces. The reason being that they do not satisfy the Sobolev-to-Lipschitz property. Moreover, the Sobolev-to-Lipschitz property is essential as the $\mathsf{BE}(K,N)$-condition alone cannot imply anything about the geometry of the space as was pointed out to us by Gigli with the following examples. Any metric space equipped with an atomic measure satisfies $\mathsf{BE}(K,N)$ for any $K$ and $N$ as the Cheeger energy is identically 0. A more geometric example is the following: For a given compact smooth embedded submanifold $M$ in $\mathbb{R}^{n}$ one can consider the induced Riemannian metric $g$ from $\mathbb{R}^{n}$ (and in turn the Riemannian distance $d_{g}$) and Riemannian volume $d\mathrm{vol}_{g}$. Then by compactness, the metric measure space $(M,d_{g},d\mathrm{vol}_{g})$ is an $\mathsf{RCD}(K,N)$-space for adequate $K$ and $N$. Now, $M$ can be endowed with the restriction of the Euclidean distance as well, i.e., $d(x,y)=\left\|x-y\right\|$ (and the same measure $d\mathrm{vol}_{g}$). It is clear that this distance in general does not make $M$ a geodesic space and thus, this metric measure structure cannot satisfy the $\mathsf{RCD}$-condition in general. However, the Cheeger energies associated to both distances agree as can be seen directly from the definition above since the Lipschitz constants of functions almost agree for both distances since they almost coincide at small scales.
Often, if there is no room for confusion, we simply refer to metric measure spaces satisfying ?THM? LABEL:D:_RCD-condition as $\mathsf{RCD}$-spaces without making any explicit mention of the parameters $K$ and $N$.
@@
Let $(X,d,\mathfrak{m})\in\mathbf{X}$ be a length space and assume that there exists a covering $\{U_{i}\}_{i\in I}$ of $X$ by non-empty open subsets such that $\mathfrak{m}(U_{i})<\infty$ if $K<0$, and $(\overline{U}_{i},d,\mathfrak{m}\llcorner\overline{U}_{i})\in X$ satisfy $\mathsf{RCD}(K,N)$. Then $(X,d,\mathfrak{m})$ is a $\mathsf{RCD}(K,N)$ space.
Let us describe some features of $\mathsf{RCD}$-spaces that will be useful to us. Their proofs can be found for example in [23]. We remark that the tensorization property in item (2) below was originally shown to hold for the $\mathsf{BE}$-condition in [6].
@@
Let $(X,d_{X},x),(Y,d_{Y},y)$ be complete and separable pointed metric spaces. Given $\epsilon>0$, we say that a map $f_{\epsilon}\colon B_{1/\epsilon}(x)\rightarrow B_{1/\epsilon}(y)$ is an $\epsilon$-Gromov-Hausdorff approximation if:
@@
    For all $u,v\in B_{1/\epsilon}(x)$, $|d_{X}(u,v)-d_{Y}(f_{\epsilon}(u),f_{\epsilon}(v))|<\epsilon$.
@@
    For all $y\in B_{1/\epsilon}(y)$, there exists some $x\in B_{1/\epsilon}(x)$ such that $d_{Y}(f_{\epsilon}(x),y)<\epsilon$.
@@
A sequence $\{(X_{n},d_{n},\mathfrak{m}_{n},x_{n})\}_{n\in\mathbb{N}}$ of complete pointed metric measure spaces converges in the pointed measured Gromov-Hausdorff sense to a complete pointed metric measure space $(Y,d_{Y},\mathfrak{m}_{Y},y)$ if it converges in the pointed Gromov-Hausdorff sense with the extra assumption that the $\epsilon_{n}$-Gromov-Hausdorff approximations are measurable and satisfy
$f_{\epsilon\#}\mathfrak{m}_{n}\rightharpoonup\mathfrak{m}_{Y}$
where the topology is the $\text{weak}^{\star}$-topology with respect to continuous functions with bounded support on $Y$.
@@
Let $(X,d,\mathfrak{m})$ be an $\mathsf{RCD}(K,N)$-space for $K\in\mathbb{R}$ and $N\in[1,\infty)$. Given a point $x_{0}\in X$, we say that a pointed metric measure space $(Y,d_{Y},\mathfrak{m}_{Y},y)$ is a tangent space of $X$ at $x_{0}$ if there exists a sequence $\{r_{i}\}_{i\in\mathbb{N}}\subset(0,1)$ such that $r_{i}\rightarrow 0$, and the sequence
@@
Given $m\in\mathbb{N}\cap[1,N]$, let $\mathcal{R}_{m}$ denote the set of points $x\in X$ such that $\mathrm{Tan}(X,d,\mathfrak{m},x)=\{(\mathbb{R}^{m},d_{\mathbb{E}},\mathcal{L}^{m},0)\}.$ Using Regular Lagrangian flows, Bruè and Semola in [13] proved that there is precisely a unique $m$ such that the set $\mathcal{R}_{m}$ has full measure.
@@ 2.3. Non-collapsed spaces
Thanks to the structure of $\mathsf{RCD}$-spaces, we can check that for non-collapsed spaces we have necessarily that $N\in\mathbb{N}$ (see Theorem $1.12$ in [22]). Non-collapsed $\mathsf{RCD}$-spaces also enjoy good structural properties which are stronger than those of more general $\mathsf{RCD}$-spaces. For example, we have the following result about the tangent spaces (see Proposition $2.8$ in [22]).
@@
Regarding the essential dimension, we have the following result by Brena, Gigli, Honda, and Zhu [12, Theorem 1.7] which lets us identify non-collapsed spaces.
@@ 2.4. Warped products
(3) $d_{X}(\gamma(t),\gamma(s))\leq\int_{s}^{t}f(r)\,dr.$
Given an AC curve $\gamma\colon[0,1]\to X$, the following limit exists for a.e. $t_{0}\in[0,1]$:
$|\dot{\gamma}(t_{0})|:=\lim_{h\to 0}\frac{d_{X}(\gamma(t_{0}+h),\gamma(t_{0}))}{|h|}.$
Moreover, the function $t\mapsto|\dot{\gamma}(t)|$ (called the metric speed of $\gamma$) is in $L^{1}(0,1)$ and is the minimal function satisfying (3).
@@
$L_{f}(\gamma):=\int_{0}^{1}\sqrt{|\dot{\alpha}(t)|^{2}+(f\circ\alpha)^{2}(t)|\dot{\beta}(t)|^{2}}\,dt.$
For convenience, we define the length of a non-admissible curve as $\infty$. $L_{f}$ is a length-structure on the class of admissible curves. We define a pseudo-metric $\tilde{d}_{f}$ on $B\times F$, between $(p,x)$ and $(q,y)$ by setting
@@
Then we have an induced equivalence relation $\sim$, where $(p,x)\sim(q,y)$ if and only if $\tilde{d}_{f}((p,x),(q,y))=0$. We define the warped product of $B$ and $F$ via $f$ as $B\times_{f}F:=B\times F/\sim$ with the distance $d_{f}([p,x],[q,y]):=\tilde{d}_{f}((p,x),(q,y))$. By definition $B\times_{f}F$ is an intrinsic metric space, and since $B$ and $F$ are complete and locally compact, then the warped product is complete and locally compact. As pointed out in [50, Remark 2.4], the warped product $B\times_{f}F$ is geodesic.
@@
$\mathrm{Con}_{K}(F)=\begin{cases}F\times[0,\pi/\sqrt{K}]/(F\times\{0,\pi/\sqrt{K}\})&\mbox{if }K>0\\ F\times[0,\infty)/(F\times\{0\})&\mbox{if }K\leq 0,\end{cases}$
@@
$:=\begin{cases}\cos^{-1}_{K}\Big(\cos_{K}(s)\cos_{K}(t)+K\sin_{K}(s)\sin_{K}(t)\cos_{K}\big(d_{F}(p,q)\wedge\pi\big)\Big)&\mbox{if }K\neq 0\\ \sqrt{s^{2}+t^{2}-2st\cos\big(d_{F}(p,q)\wedge\pi\big)}&\mbox{if }K=0,\end{cases}$
@@
Here, for $I_{K}=[0,\pi/\sqrt{K}]$ for $K>0$ and $0,\infty)$ for $K\leq 0$, we have $\cos_{K}\colon I_{K}\to[0,\infty)$ given by
$\cos_{K}(t):=\begin{cases}\cos\left(\sqrt{K}t\right)&\mbox{if }K>0\\ \cosh\left(\sqrt{-K}t\right)&\mbox{if }K<0,\end{cases}$
@@
[…]
@@
If $\mathrm{diam}(F)\leq\pi$ and $F$ is a length space, then the $(K,N)$-cone $\mathrm{Con}^{N}_{K}(F)$ coincides with the $N$-warped product $I_{K}\times_{\sin_{K}}^{N}F$ (e.g. [14, Theorem 3.6.17]).